# Highly Efficient Superconducting Diodes and Rectifiers for Quantum Circuitry

### v1

@@ Abstract
Figure 1: Efficient and reproducible superconducting diode (SD) effect in V/EuS devices. (a) The left panel is an optical microscope image of an SD where the dark areas have been etched. The scale bar is 20 $\mu$m. The right panel is a schematic of the superconducting rectifier with the same colors as the left image. Red rectangles highlight the SDs, and the outer square size is $1\times 1$ mm2. (b) $I$-$V$ characteristics of a single SD for $B_{z}=-5$ Oe and $M_{\mathrm{EuS}}$ along $-y$. The arrows indicate the sweep directions and the forward, reverse, and retrapping currents ($I_{\mathrm{c}}^{+}$, $I_{\mathrm{c}}^{-}$, and $I_{\mathrm{r}}$, respectively) are defined. The upper left inset illustrates the SD induced by stray magnetic fields from the EuS (green) acting on the V (orange). The lower right inset shows the SD effect induced by the out-of-plane magnetic field ($B_{z}$) in combination with edge asymmetry. (c) and (d) $I_{\mathrm{c}}^{+}$ and $I_{\mathrm{c}}^{-}$ of the SD in panel b as a function of $B_{z}$ after aligning the EuS magnetization as indicated by the insets. The black dashed lines are fits to the model described in the text. The remanence of the setup magnet is corrected assuming that the $I_{\mathrm{c}}^{+}$ and $I_{\mathrm{c}}^{-}$ peaks are centered around zero. (e) Schematic representation of a superconducting diode bridge with the circuit used to measure the SD effect. The SDs are sketched as blue diodes. (f) SD effect in a superconducting diode bridge at $B_{z}\approx\pm 4$ Oe. The measurement temperature is 1.7 K.
@@
Figure 2: Reversible operation and efficiency of a superconducting rectifier. (a) Schematic of the measurement circuit used to measure the rectification of a superconducting diode bridge rectifier. $R$ is the resistance of the output load. (b) Experimental demonstration of reversible superconducting diode operation for $R=2\,\Omega$. The upper panel shows the input current ($I_{\mathrm{ac}}$), measured for improved accuracy and smoothed using a moving average filter with a 7-point window. The middle and lower panels show the output voltages at $B_{z}=\pm 4$ Oe, respectively. The yellow stripes highlight the time intervals where the reverse diodes are in the normal state ($|I_{\mathrm{ac}}{}|\geq 2\min(I_{\mathrm{c}}^{+}{},\,I_{\mathrm{c}}^{-}{})$). The $V_{\mathrm{out}}$ amplitude is slightly lower in the lower panel due to a $\sim 5$% lower $I_{\mathrm{ac}}$ (not shown for simplicity), an offset of 0.1 mV has been corrected from $V_{\mathrm{out}}$, and the measurement temperature is 1.7 K. (c) Electrical circuit used to simulate the rectifier performance. The input current ($I_{\mathrm{in}}$) is constant and sufficiently large to turn SD2 and SD3 resistive (see labels), acquiring a resistance $R_{\mathrm{n}}$ while SD1 and SD4 remain superconducting. (d) Maximal output signal as a function of $R$. The dots are experimental values, and the lines are predictions of the resistor model in panel c for different $R_{\mathrm{n}}$. For $10<R<200\,\Omega$, $V_{\mathrm{out}}$ peaks with two different amplitudes are observed. Their measurement probability ($P$) determines the color of the points according to the color bar. (e) Equivalent resistor circuit to panel c. (f) Same as panel e but considering that SD1 or SD4 become resistive. (g) Rectifier efficiency vs. $R$ extracted from panel c. The black and red lines are obtained from the models in panels e and f, respectively with $R_{\mathrm{n}}{}=158\,\Omega$.
@@
SDs are systems where inversion and time reversal symmetries are broken [11], a condition that can be achieved using different approaches. External magnetic fields [12, 13, 14, 15, 16], ferromagnets [17, 18, 19, 15], and trapping Abrikosov vortices in Josephson junctions [20] can break time-reversal symmetry. Moreover, inversion symmetry can be broken using materials with low crystallographic symmetry [21, 17, 22], introducing device-scale asymmetries [23, 12, 24, 25, 26, 14, 15], or even applying inhomogeneous magnetic fields [27, 23]. Thanks to this versatility, a wide variety of approaches have been employed to realize SDs [12, 13, 14, 15, 21, 18, 19, 20, 21, 17, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 16, 39], opening the door for potential applications.
@@
which is shown by the solid lines in Fig. 2d. Using Eq. 1 to fit the measured data by least squares gives $R_{n}=158\,\Omega$, which is comparable with the $R_{n}\approx 200\,\Omega$ obtained from the slope of the $I$-$V$ curves in Fig. 1f. We attribute this slight discrepancy to small differences in $R_{n}$ between the SDs. The current efficiency is represented in Fig. 2d, where the black line is the expected efficiency
@@
extracted from Fig. 2e with $R_{n}=158\,\Omega$, where $I_{\mathrm{out(in)}}$ and $P_{\mathrm{out(in)}}$ are the output (input) current and power, respectively. The black dots are obtained from the experimental data and agree well with the model for high $R$. However, for $R<200\,\Omega$, the data in Fig. 2e shows two sets of peaks with a factor of two difference in the amplitude (Fig. S3) which do not agree with Eq. 2. The dot color indicates the probability ($P$) by which they occur in the measurements according to the color bar in Fig. 2d. We explain these lower-amplitude peaks considering that either SD1 or SD4 becomes normal. This results in the equivalent circuit shown in Fig. 2f, which gives the following expression for the current efficiency
@@
When $R\ll R_{\mathrm{n}}{}$, the efficiency given by Eq. 3 is two times smaller than Eq. 2. This is caused by the increase in the resistance of the current path that goes through the load resistor $R$, which penalizes the circuit efficiency even for $R=0$. Using $R_{n}=158\,\Omega$, Eq. 3 results in the red line in Fig. 2d, which shows good agreement with the experiment at low $R$. This result is independent of which SD is turning normal (SD1 or SD4). Two possible reasons can explain this premature transition to the resistive state: 1. Heating induced by SD2 and SD3 decreases $I_{\mathrm{c}}^{+}$ of SD1 and SD4. 2. Small differences in $I_{\mathrm{c}}^{+}$ between the SDs in the bridge. These effects do not occur for higher $R$ because the current load on the forward diodes decreases with increasing $R$. Nevertheless, the maximal efficiency achieved is 43%, a remarkable metric for such a new device.
@@
To conclude, by combining type A and type C SD effects in V/EuS thin films, we have demonstrated robust single diodes with enhanced efficiency and tunability. With the superb efficiency and reproducibility achieved, we constructed the first superconducting rectifiers operating with frequencies up to 40 kHz and efficiencies up to 43%. The integration of SC rectifiers in SC electronic chips is a key step towards the scaling of ERSFQ-based quantum and classical computers [3], which would also benefit new sensing applications [4]. In addition, DC power delivery is not the only problem that SD circuits can tackle. Circulators are ubiquitous devices in the majority of quantum computing systems. The existing devices are bulky, expensive 3D objects increasing quantum system complexity and cost. Their integration in SD promises to boost quantum computation [41].
@@ I Acknowledgements
We acknowledge Dr. A. Kirichenko for insightful discussions. This work was supported by Air Force Office of Sponsored Research (FA9550-23-1-0004 DEF), Office of Naval Research (N00014-20-1-2306), National Science Foundation (NSF-DMR 2218550 and 1231319); Army Research Office (W911NF-20-2-0061, DURIP W911NF-20-1-0074). EC and PW acknowledge the NSF CAREER DMR-2046648.

### v2

@@ Abstract
Figure 1: Efficient and reproducible superconducting diode (SD) effect in V/EuS devices. (a) The left panel is an optical microscope image of an SD where the dark areas have been etched. The scale bar is 20 $\mu$m. The right panel is a schematic of the superconducting rectifier with the same colors as the left image. Red rectangles highlight the SDs, and the outer square size is $1\times 1$ mm2. (b) $I$-$V$ characteristics of a single SD for $B_{z}=-5$ Oe and $M_{\mathrm{EuS}}$ along $-y$. The arrows indicate the sweep directions and the forward, reverse, and retrapping currents ($I_{\mathrm{c}}^{+}$, $I_{\mathrm{c}}^{-}$, and $I_{\mathrm{r}}$, respectively) are defined. The upper left inset illustrates the SD induced by stray magnetic fields from the EuS (green) acting on the V (orange). The lower right inset shows the SD effect induced by the out-of-plane magnetic field ($B_{z}$) in combination with edge asymmetry. (c) and (d) $I_{\mathrm{c}}^{+}$ and $I_{\mathrm{c}}^{-}$ of the SD in panel b as a function of $B_{z}$ after aligning the EuS magnetization as indicated by the insets. The black dashed lines are fits to the model described in the text. The remanence of the setup magnet is corrected assuming that the $I_{\mathrm{c}}^{+}$ and $I_{\mathrm{c}}^{-}$ peaks are centered around zero. The error bars corresponding to one standard deviation are smaller or equal to the dot size. (e) Schematic representation of a superconducting diode bridge with the circuit used to measure the SD effect. The SDs are sketched as blue diodes. (f) SD effect in a superconducting diode bridge at $B_{z}\approx\pm 4$ Oe. The measurement temperature is 1.7 K.
@@
Figure 2: Reversible operation and efficiency of a superconducting rectifier. (a) Schematic of the measurement circuit used to measure the rectification of a superconducting diode bridge rectifier. $R$ is the resistance of the output load. (b) Experimental demonstration of reversible superconducting diode operation for $R=2\,\Omega$. The upper panel shows the input current ($I_{\mathrm{ac}}$), measured for improved accuracy and smoothed using a moving average filter with a 7-point window. The middle and lower panels show the output voltages at $B_{z}=\pm 4$ Oe, respectively. The yellow stripes highlight the time intervals where the reverse diodes are in the normal state ($|I_{\mathrm{ac}}{}|\geq 2\min(I_{\mathrm{c}}^{+}{},\,I_{\mathrm{c}}^{-}{})$). The $V_{\mathrm{out}}$ amplitude is slightly lower in the lower panel due to a $\sim 5$% lower $I_{\mathrm{ac}}$ (not shown for simplicity), an offset of 0.1 mV has been corrected from $V_{\mathrm{out}}$, and the measurement temperature is 1.7 K. (c) Electrical circuit used to simulate the rectifier performance. The input current ($I_{\mathrm{in}}$) is constant and sufficiently large to turn SD2 and SD3 resistive (see labels), acquiring a resistance $R_{\mathrm{n}}$ while SD1 and SD4 remain superconducting. (d) Maximal output signal as a function of $R$. The dots are experimental values, and the lines are predictions of the resistor model in panel c for different $R_{\mathrm{n}}$. For $10<R<200\,\Omega$, $V_{\mathrm{out}}$ peaks with two different amplitudes are observed. Their measurement probability ($P$) determines the color of the points according to the color bar. (e) Equivalent resistor circuit to panel c. (f) Same as panel e but considering that SD1 or SD4 become resistive. (g) Rectifier efficiency vs. $R$ extracted from panel c. The black and red lines are obtained from the models in panels e and f, respectively with $R_{\mathrm{n}}{}=154\,\Omega$. The dots in panels d and g are larger than the error bars corresponding to one standard deviation and the uncertainty of $R_{\mathrm{n}}$ in panel d corresponds to the fitting error.
@@
SDs are systems where inversion and time reversal symmetries are broken [11], a condition that can be achieved using different approaches. External magnetic fields [12, 13, 14, 15, 16], ferromagnets [17, 18, 19, 15], and trapping Abrikosov vortices in Josephson junctions [20] can break time-reversal symmetry. Moreover, inversion symmetry can be broken using materials with low crystallographic symmetry [21, 17, 22], introducing device-scale asymmetries [23, 12, 24, 25, 26, 14, 15], or even applying inhomogeneous magnetic fields [27, 23]. Thanks to this versatility, a wide variety of approaches have been employed to realize SDs [12, 13, 14, 15, 21, 18, 19, 20, 21, 17, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 16, 39, 40, 41, 42, 43, 44], opening the door for potential applications.
@@
which is shown by the solid lines in Fig. 2d. Using Eq. 1 to fit the measured data by least squares gives $R_{n}=156\,\Omega$, which is comparable with the $R_{n}\approx 200\,\Omega$ obtained from the slope of the $I$-$V$ curves in Fig. 1f. We attribute this slight discrepancy to small differences in $R_{n}$ between the SDs. The current efficiency is represented in Fig. 2d, where the black line is the expected efficiency
@@
extracted from Fig. 2e with $R_{n}=156\,\Omega$, where $I_{\mathrm{out(in)}}$ and $P_{\mathrm{out(in)}}$ are the output (input) current and power, respectively. The black dots are obtained from the experimental data and agree well with the model for high $R$. However, for $R<200\,\Omega$, the data in Fig. 2e shows two sets of peaks with a factor of two difference in the amplitude (Fig. S3) which do not agree with Eq. 2. The dot color indicates the probability ($P$) by which they occur in the measurements according to the color bar in Fig. 2d. We explain these lower-amplitude peaks considering that either SD1 or SD4 becomes normal. This results in the equivalent circuit shown in Fig. 2f, which gives the following expression for the current efficiency
@@
When $R\ll R_{\mathrm{n}}{}$, the efficiency given by Eq. 3 is two times smaller than Eq. 2. This is caused by the increase in the resistance of the current path that goes through the load resistor $R$, which penalizes the circuit efficiency even for $R=0$. Using $R_{n}=156\,\Omega$, Eq. 3 results in the red line in Fig. 2d, which shows good agreement with the experiment at low $R$. This result is independent of which SD is turning normal (SD1 or SD4). Two possible reasons can explain this premature transition to the resistive state: 1. Heating induced by SD2 and SD3 decreases $I_{\mathrm{c}}^{+}$ of SD1 and SD4. 2. Small differences in $I_{\mathrm{c}}^{+}$ between the SDs in the bridge. These effects do not occur for higher $R$ because the current load on the forward diodes decreases with increasing $R$. Nevertheless, the maximal efficiency achieved is 42$\pm$2%, a remarkable metric for such a new device.
@@
To conclude, by combining type A and type C SD effects in V/EuS thin films, we have demonstrated robust single diodes with enhanced efficiency and tunability. With the superb efficiency and reproducibility achieved, we constructed the first superconducting rectifiers operating with frequencies up to 40 kHz and efficiencies up to 42$\pm$2%. The integration of SC rectifiers in SC electronic chips is a key step towards the scaling of ERSFQ-based quantum and classical computers [3], which would also benefit new sensing applications [4]. In addition, DC power delivery is not the only problem that SD circuits can tackle. Circulators are ubiquitous devices in the majority of quantum computing systems. The existing devices are bulky, expensive 3D objects increasing quantum system complexity and cost. Their integration in SD promises to boost quantum computation [46].
@@ I Acknowledgements
We acknowledge Dr. M. B. Ketchen, Dr. A. Kirichenko, and A. Gupta for insightful discussions and M. Mondol for technical assistance. This work was supported by Air Force Office of Sponsored Research (FA9550-23-1-0004 DEF), Office of Naval Research (N00014-20-1-2306), National Science Foundation (NSF-DMR 2218550 and 1231319); Army Research Office (W911NF-20-2-0061, DURIP W911NF-20-1-0074). EC and PW acknowledge the NSF CAREER DMR-2046648. This work was carried out in part through the use of MIT.nano’s facilities.